Fermi National Laboratory
 

Accelerator Update for March 16, 2007 - March 19, 2007

 

Friday March 16

The day shift began with Operations monitoring the Tevatron store 5284 (156.61E30) and the Recycler stash (3.6E10), with the Antiproton Source (Pbar) stacking (59.5E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI and the Meson MTest experiment taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

Pbar experts began a reverse proton study at 12:32 PM.

Operations prepared the NuMI MI-65 area for an access at 1:12 PM.

A Recycler expert began a study at 1:27 PM.

An MI expert began a one-hour mixed-mode study at 1:34 PM.

NuMI experts reported at 2:50 PM that they found an LCW leak on V118-4. LCW technicians began the repair.

Operations reported at 3:59 PM that the Recycler had stashed antiprotons seven times during the day shift.

Pbar experts completed their study. Stacking resumed at 4:41 PM.

The NuMI MI-65 LCW leak was repaired by 4:49 PM. NuMI resumed taking beam at 5:03 PM.

At 5:36 PM, a Linac RF station (LRF4) tripped off for the first of a number of times in one hour. A Linac expert began repairs. Later, other experts came in to help with the repair.

The MTest experiment reported at 6 PM that they were finished with the beam for the night.

Linac experts brought LRF4 online at 8:49 PM. Stacking resumed and then LRF4 tripped off again five minutes later.

LRF4 came back on line at 11:50 PM. Pbar resumed stacking and NuMI resumed taking beam.

At 11:59 PM, Operations reported that the lithium lens was pulsing intermittently. Operators reset the device.

Saturday March 17

The midnight shift began with Operations monitoring store 5284 (50.97E30) and the stash (78.5E10), with Pbar stacking (39.3E10), and with NuMI taking beam.

At 12:01 AM, the NuMI beam permit tripped off. It reset.

At 2:53 AM, a Booster RF station (BRF6) tripped off. It reset.

Operations established beam to MTest at 6:34 AM.

At 7:43 AM, Operations reported that the H- Source began sparking so badly that it wasn’t putting out much beam. Operators accessed the Preaccelerator dome to reset a device.

Operations reported at 7:59 AM that the Recycler stashed three times during the midnight shift.

Linac beam resumed at 8:20 AM, but Operations reported that the H- Source wasn’t running well. A Linac expert arrived and began baby-sitting the source around 8:30 AM.

At 11:27 AM, BRF6 tripped off. It reset.

Operations terminated store 5284 at noon.

Operations prepared D0 for an access at 12:10 PM. They came out of access ninety minutes later.

Operations put the TeV into a dry squeeze at 2:01 PM.

TeV experts began a study period at 2:27 PM.

Operations put the TeV through a dry squeeze at 5:02 PM.

Operations put the TeV through a wet squeeze at 5:52 PM.

Operations put the TeV into shot setup at 6:13 PM.

Operations established store 5289 at 7:43 PM with an initial luminosity of 205.38E30.

Operations reported that the Recycler had stashed four times during the last twelve hours.

Operations reported at 8 PM that store 5298’s luminosity was 135.69E30, that the stack was at 59.7E10, and the stash was 21.2E10.

At 8:38 PM, Operations reported that they had been having problems with the TeV beam’s halo. Operators contacted and expert.

Pbar experts began a stacktail study at 9:15 PM.

At 9:19 PM, Operations reported that they had found a collimator that hadn’t gone to its correct position. Operators moved it to the correct position.

Pbar experts completed their study. Pbar resumed stacking at 10:03 PM.

At 11:05 PM, BRF6 tripped off. It reset.

Sunday March 18

The midnight shift began with Operations monitoring store 5289 and the stash, with Pbar stacking, and with NuMI taking beam.

At 1:59 AM, Operations reported that the Pbar target had stopped rotating. Operators switched the rotation from reverse to normal and the target resumed rotating.

At 2:13 AM, the Pbar target stopped rotating again. Operations talked with experts who said that the target area would have to be accessed. Operations put Pbar into standby.

Work began on the Pbar target around 6:32 AM. The work ended about an hour later. The Pbar expert reported that he had replaced the target’s motor.

Pbar resumed stacking at 7:49 AM.

Operations reported at 8:04 AM that the Recycler stashed three times during the last twelve hours.

At 8:54 AM, Operations reported that a TeV RF station (TRF8) had tripped off.

At 8:58 AM, Operation reported that they found several MI vacuum pumps off. Operators reset the pumps.

At 9:04 AM, Operations reported that the Booster vacuum valves closed when the West Anode Power Supply tripped off. Operators reset the power supply and everything else.

MTest resumed taking beam at 9:29 AM.

Operations reported at 7:55 PM that the Recycler had stashed four times during the last twelve hours.

Operations reported at 8 PM that store 5298’s luminosity was 34.06E30, the stack was 60.0E10, and the stash was 284.0E10.

Operations terminated store 5289 at 9:04 PM.

Operators reset TRF8 (with the help of an expert) at 9:18 PM.

Operations put the TeV into shot setup at 9:19 PM.

At 11:54 PM, Operations reported that the protons and antiprotons had been loaded for the next store, but that tune problems were preventing Operations from calling it a store.

Monday March 19

The midnight shift began with operators trying to resolve TeV tune problems, with Operations monitoring the stash, with Pbar stacking, and with NuMI taking beam.

Operations established store 5290 at 12:48 AM with an initial luminosity of 182.77E30.

At 4:27 AM, BRF6 tripped off. It reset.

MTest resumed taking beam at 6:41 AM.

Operations reported at 7:59 AM that the Recycler had stashed 5 times during the last twelve hours.

Pbar began an eight-hour stacktail equalizer study.

The Plans for Monday

The plan for today is to keep store 5290 until sometime during the midnight shift on Tuesday. Pbar will conduct an eight-hour stacktail equalizer study during the day shift; its newly installed equalizer will, hopefully, increase the stacking rate. Recycler experts will also conduct a study.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/19/2007    email Fermilab